OpenStudy (anonymous):

Flora has $50 to spend on clothes. She wants to buy a pair of jeans for $20 and spend the rest on t-shirts. Each t-shirt costs $6. Which inequality represents the number of t-shirts she can purchase? x ≥ 5 x ≤ 5 x ≤ 11 x ≥ 11

OpenStudy (anonymous):

@Abhisar

OpenStudy (abhisar):

What do you think it should be?

OpenStudy (anonymous):

c?

OpenStudy (abhisar):

She has $50 in total and she is gonna spend $20 on t-shirt. How much money left?

OpenStudy (anonymous):

30

OpenStudy (abhisar):

Right! Since, each jeans cost $6, what is the maximum number she can buy with the left $30?

OpenStudy (anonymous):

5 pairs of jeans

OpenStudy (abhisar):

Right! c: So she can buy 5 or less than 5 jeans. What should be the correct inequality representation?

OpenStudy (anonymous):

b

OpenStudy (abhisar):

YaY !!

OpenStudy (anonymous):

The sum of two numbers is 60. The greater number is 6 more than the smaller number. Which equation can be used to solve for the smaller number? x + (x + 6) = 60 x − (x + 6) = 60 x(x + 6) = 60 x(x− 6) = 60

OpenStudy (anonymous):

@Abhisar